The company that operates for 18 hours a day has introduced a new daily wage system that
incorporates variable hourly rate to encourage longer shifts. The company pays $10 for first
hour or any part of the first hour. Afterwards, company pays additional 2% per hour for each
hour or part of each additional hour. Write a piece-wise function that shows the salary as a
function of hours. Also sketch a graph of this function.
1
Expert's answer
2021-09-26T17:46:34-0400

Let a person work h hours for the company

Note if "0<h\le 1" then the company pays $10

If "1<h\le2" then the company pays $"10+10+\frac{2*10}{100}"

If "2<h\le 3" then the company pays $"10+10+10+\frac{2*(3-1)*10}{100}"

="30+\frac{2*(3-1)*10}{100}"

So for h hours the company pays "10*h+\frac{2*(h-1)*10}{100}" ,h>0

So the function is;

"\begin{cases}
 10 &\text{if } 0<h\le1 \\
 10+10+\frac{1}{5} &\text{if } 1<h\le 2 \\
10+10+10+\frac{2}{5} &\text{if } 2<h\le 3 \\10+10+10+10+\frac{3}{5} &\text{if }3<h\le 4
\end{cases}"

The graph is as below,